12. Transactions for the month of June were: June 1 Purchases (balance) 2,200 @ $3.20 3,700 @ 3.10 3,400 @ 3.30 2,600 @ 3.40Transactions for the month of June were:PurchasesSales (units)June 1(balance) 2,200 @$3.20June 21,200 33,700 @3.1063,200 73,400 @3.3092,000 152,600 @3.4010800 221,800 @3.50182,800 25400 Assuming that perpetual inventory records are kept in units only, the cost of goods sold on a LIFO basis is
Answer #1

Calculate Cost of goods sold Under LIFO basis

June 2 1200*3.20 = 3840
June 6 3200*3.1 = 9920
June 9 2000*3.3 = 6600
June 10 800*3.3 = 2640
June 18 2600*3.4+200*3.3 = 9500
June 25 400*3.5 = 1400
Total Cost of goods sold 33900

So answer is $33900